# Style over Substance: Distilled Language Models Reason Via Stylistic Replication

### v1

@@ Abstract
Specialized reasoning language models (RLMs) have demonstrated that scaling test-time computation through detailed reasoning traces significantly enhances performance. Although these traces effectively facilitate knowledge distillation into smaller, instruction-tuned models, the precise nature of transferred reasoning remains unclear. In this study, we investigate to what extent distilled models internalize replicated stylistic patterns during reasoning. To this end, we systematically analyze reasoning traces, identifying structural and lexical patterns that characterize successful reasoning. We then introduce two new datasets – a dataset of emergent reasoning traces and a synthetic dataset explicitly constructed to replicate these stylistic patterns – to precisely examine their influence on distilled models’ reasoning capabilities. We find that models trained on the synthetic traces achieve comparable performance, indicating that distilled reasoning abilities rely significantly on surface-level patterns. Surprisingly, we observe an increase in performance even when the synthetic traces are altered to lead to the wrong answer. Our findings highlight how stylistic patterns can be leveraged to efficiently enhance LM reasoning across diverse model families.
@@ 3.3 Experimental details
Baselines To contextualize our findings, we evaluate (1) the base instruction-tuned model, (2) the model after SFT on ST, and (3) the model after SFT on ST-HC. This comprehensive evaluation allows us to isolate the impact of our hard-coded reasoning approach across different model sizes and compare it against both emergent reasoning in RLMs and reasoning in instruction-tuned LMs. For completeness, we evaluate the teacher models used to generate our datasets as well.
@@ 4.1 Hard-coded reasoning results
The performance across all combinations of models and datasets evaluated in our study is presented in table 1. Our results clearly indicate that models finetuned with structured reasoning traces, whether emergent (ST) or synthetic (ST-HC), consistently and significantly outperform their baseline instruction-tuned counterparts across all benchmarks. Notably, even the smaller 3B and 8B parameter models exhibit substantial performance improvements. For instance, the Llama 3.2 3B model gains over 31 absolute percentage points on MATH500 when finetuned on ST, and the Ministral 8B model sees its AIME2024 score more than triple with either ST or ST-HC finetuning. These findings demonstrate that smaller models can markedly benefit from structured reasoning finetuning, challenging previous suggestions of minimal improvements for models of this scale (Team, 2025). Comparing the two fine-tuning approaches, we observe that models trained on synthetic traces achieve performance that approaches, and in some cases matches those trained on emergent RLM traces. This demonstrates that replicating the style of reasoning, even using a less capable teacher model guided by our prompt, is highly effective for distilling reasoning capabilities.
@@
Table 1: Performance comparison of language models finetuned on reasoning traces. We compare base models against versions finetuned on ST (emergent traces from R1) and ST-HC (synthetic traces from GPT-4o with the prompt in figure 3). This evaluates the impact of reasoning trace style on downstream performance across model scales and benchmarks. GPT-4o HC refers to the base LM prompted with our structure from figure 3. All results are pass@1 accuracy as a percentage.

### v3

@@ Abstract
Specialized reasoning language models (RLMs) have demonstrated that scaling test-time computation through detailed reasoning traces significantly enhances performance. Although these traces effectively facilitate knowledge distillation into smaller, instruction-tuned models, the precise nature of transferred reasoning remains unclear. In this study, we investigate to what extent distilled models internalize replicated stylistic patterns during reasoning. To this end, we systematically analyze reasoning traces, identifying structural and lexical patterns that characterize successful reasoning. We then introduce two new datasets – a dataset of emergent reasoning traces and a synthetic dataset explicitly constructed to replicate these stylistic patterns – to precisely examine their influence on distilled models’ reasoning capabilities. We find that models trained on the synthetic traces achieve comparable performance, indicating that distilled reasoning abilities rely significantly on surface-level patterns. Surprisingly, we observe an increase in performance even when the synthetic traces are altered to lead to the wrong answer. Our findings highlight how stylistic patterns can be leveraged to enhance LM reasoning across model families.
@@ 3.3 Experimental details
Baselines To contextualize our findings, we evaluate (1) the base instruction-tuned model, (2) the model after SFT on ST, and (4) the model after SFT on ST-HC. Additionally, to gauge the impact of the particular style replication we propose versus distilling from regular CoT, we add another baseline, where we instruct the generating LLM to think step-by-step (SBS), following Kojima et al. (2023). We then use the resulting CoT for distillation – similarly to how we use the emergent (ST) and synthetic (ST-HC) reasoning traces. This comprehensive evaluation allows us to isolate the impact of our hard-coded reasoning approach across different model sizes and compare it against the reasoning capabilities of the model itself. For completeness, we evaluate the teacher models used to generate our datasets as well.
@@ 4.1 Hard-coded reasoning results
The performance across all combinations of models and datasets evaluated in our study is presented in table 1. Our results clearly indicate that models finetuned with structured reasoning traces, whether emergent (ST) or synthetic (ST-HC), consistently and significantly outperform their baseline instruction-tuned counterparts across all benchmarks. Notably, even the smaller 3B and 8B parameter models exhibit substantial performance improvements. For instance, the Llama 3.2 3B model gains over 31 absolute percentage points on MATH500 when finetuned on ST, and the Ministral 8B model sees its AIME2024 score more than triple with either ST or ST-HC finetuning. These findings demonstrate that smaller models can markedly benefit from structured reasoning finetuning, challenging previous suggestions of minimal improvements for models of this scale (Team, 2025). Comparing the two fine-tuning approaches, we observe that models trained on synthetic traces achieve performance that approaches, and in some cases matches those trained on emergent RLM traces. This demonstrates that replicating the style of reasoning, even using a less capable teacher model guided by our prompt, is highly effective for distilling reasoning capabilities. Finally, the results show that while generating synthetic data with a SBS prompt is beneficial compared to the base model, there is a significant performance gap between SBS and our ST and ST-HC methods. This demonstrates that the performance improvements are indeed substantially driven by the specific stylistic patterns we identified, rather than solely by distilling the generating LLM’s general reasoning abilities.
@@
Table 1: Performance comparison of language models finetuned on reasoning traces. We compare base models against versions finetuned on SBS (step-by-step thinking GPT-4o), ST (emergent traces from R1), and ST-HC (synthetic traces from GPT-4o with the prompt in figure 3). This evaluates the impact of reasoning trace style on downstream performance across model scales and benchmarks. GPT-4o HC refers to the base LM prompted with our structure from figure 3. All results are pass@1 accuracy as a percentage.